At the beginning of an experiment, there are 400 grams of contaminants. Each hour, a fourth of the contaminants are filtered out.
A.Formulate a recursive sequence modeling the number of grams after n hours.
B.Use the model to calculate the amount of contaminants after the third hour of the experiment.

Answer by stanbon(75887)

At the beginning of an experiment, there are 400 grams of contaminants. Each hour, a fourth of the contaminants are filtered out.
A.Formulate a recursive sequence modeling the number of grams after n hours.
a(n) = 0.75*a(n-1)
------------------------
B.Use the model to calculate the amount of contaminants after the third hour of the experiment.
Ans: 400 ; (3/4)400 = 300 ; (3/4)300 = 900/4 ; (3/4)900/4 = 2700/16 = 168.75g's
-------------------
Cheers,
Stan H.
----------------------